# Proximal Interacting Particle Langevin Algorithms

### v2

@@ Results.
The evolution of the relative error of the missing entries of the tensor $(Y^{1},\dots,Y^{L})$ across steps is shown in Figure 7. Figure 8 displays the errors of the entries of $Y^{l}$ for a randomly chosen $l$ and a randomly chosen particle of the final cloud.
Figure 7: Relative error of the missing entries of the tensor $(Y^{1},\dots,Y^{L})$ across steps.
Figure 8: Errors of the entries of $Y^{l}$ for a randomly chosen $l$ and a randomly chosen particle of the final cloud.
D.4.2 Example 2
Dataset.
The entries of $\Phi$ are obtained from a $N(\theta,\sigma^{2})$, where $\sigma=0.25$ and $\theta=1.4$ for the experiment in the main text. The matrix $X$ is generated in the same way as the $X^{l}$ matrices in Example 1. Similarly, each entry of $Y$ is observed with a probability $0.5$.
Proximal operator of $g_{2}$.
Recall that $g_{2}(\xi,\theta,X,\Phi)$ is of the form
$g_{2}(\xi,\theta,X,\Phi)=-\log(C(\xi))+e^{-\xi}\|X\|_{\text{tr}}.$
As in Example 1, due to the difficulty of calculating $C(\xi)$ we assume that it is approximately constant $C(\xi)\approx C$. Therefore, it is not necessary to take it into account when deriving the proximal operator, which is given by
$\operatorname{prox}_{g_{2}}^{\lambda}(\xi,\theta,X,\Phi)=\operatorname*{arg\,min}_{(\alpha,\delta,Z,\Gamma)}\;\{e^{-\alpha}\|Z\|_{\text{tr}}+\frac{1}{2\lambda}\big{(}\|\xi-\alpha\|^{2}+\|\theta-\delta\|^{2}+\|X-Z\|_{F}^{2}+\|\Phi-\Gamma\|_{F}^{2}\big{)}\}.$
The minimisers $(\alpha,\delta,Z,\Gamma)$ satisfy the following system of equations
$\delta=\theta,\;\Gamma=\Phi,$
$\alpha=\xi+\lambda e^{-\alpha}\|S_{e^{-\alpha}\lambda}(X)\|_{\text{tr}}\Longrightarrow(\alpha-\xi)e^{\alpha-\xi}=\lambda e^{-\xi}\|S_{e^{-\alpha}\lambda}(X)\|_{\text{tr}},$ (58)
$Z=S_{e^{-\alpha}\lambda}(X).$ (59)
To avoid the computational cost of solving (58)-(59) using an iterative solver, we approximate (58) by
$(\alpha-\xi)e^{\alpha-\xi}\approx\lambda e^{-\xi}\|S_{e^{-\xi}\lambda}(X)\|_{\text{tr}}\Longrightarrow\alpha\approx\xi+W(\lambda e^{-\xi}\|S_{e^{-\xi}\lambda}(X)\|_{\text{tr}}).$
Substituting this $\alpha$ in (59), we obtain
$Z\approx S_{e^{-\alpha}\lambda}(X).$
Implementation.
To stabilise the implementation of MYIPLA, we have divided the gradients and proximal mapping term of the updates of $\xi$ and $\theta$ by the dimension of the tensor $X$ ($d_{x}=r\times n$) and $\Phi$ ($d_{\Phi}=m\times r$), respectively. We then set $\gamma=0.05$ and $\lambda=0.35$.
Results.
The relative error of the missing entries of the matrix $Y$ and the average relative error of the particle cloud $\Phi_{l}^{1},\dots,\Phi_{l}^{N}$ are plotted across steps in Figure 9.
Figure 9: Performance measures across steps.

### v3

@@ Results.
Figure 13 displays the original and observed matrices alongside the reconstructed matrices obtained using our different proximal algorithms. The methods are run for 3000 iterations (with a burn-in of 100 iterations) and $N=10$ particles. The NMSEs for the entire matrix and the missing entries for the final particle cloud are displayed in Table 11, together with the computation times.
(a) Original
(b) Observed
(c) MYPGD
(d) MYIPLA
(e) PIPGLA
Figure 13: Low-rank matrix completion. All the algorithms use $N=10$ particles and are run for 3000 iterations. The blue pixels in (b) represent the mask.
Table 11: Low-rank matrix completion. Normalised mean squared errors (NMSE) for the entire matrix and the missing entries achieved using the final particle cloud with $N=10$ and $3000$ iterations.
[…]
MYPGD $1.21\pm 0.49$ $1.67\pm 0.52$ $\mathbf{4.1}$
MYIPLA $\mathbf{1.13\pm 0.48}$ $\mathbf{1.53\pm 0.44}$ $4.7$
PIPGLA $2.02\pm 0.29$ $2.11\pm 0.31$ $5.5$
E.6 Ablation Study
In this section, we analyse how the choice of the regularisation parameter $\lambda$ in the Moreau–Yosida approximation affects the performance and stability of the algorithm.
Choosing an appropriate value for $\lambda$ is a challenging task as this parameter controls both the level of regularisation and the closeness to the target, and is closely tied to the step size parameter $\gamma$. Durmus et al. [2018] provides some empirical guidance on the choice of $\gamma,\lambda$ for sampling tasks. Crucinio et al. [2025] shows that $\lambda\leq\gamma$ generally leads to better results in the case of grad Lipschitz potentials, while one should choose $\lambda\geq\gamma$ for light tail distributions. Adaptive strategies to choose $\lambda$ have been considered in the optimisation literature (see Oikonomidis et al. [2024] and references therein) but equivalent results for sampling have not been obtained yet.
We conduct additional experiments to analyse the impact of the regularisation parameter $\lambda$ in the Bayesian logistic regression task with Laplace prior. In Figure 14, we report the performance (measured by NMSE) of MYIPLA, MYPGD and PIPGLA algorithms using approximate proximity maps, evaluated over a fine grid of $\lambda$ values. The step size parameters used are those listed in Table 5: $\gamma=0.05$ for MYIPLA and MYPGD, and $\gamma=0.01$ for PIPGLA. Each configuration is run with 100 different random seeds to compute confidence intervals. We observe that our algorithms exhibit stable performance across a broad range of $\lambda$ values.
(a) MYPGD, $\gamma=0.05$
(b) MYIPLA, $\gamma=0.05$
(c) PIPGLA, $\gamma=0.01$
Figure 14: Normalised MSE (%) for different values of the regularisation parameter $\lambda$ and a fixed step size $\gamma$. Each configuration is run with 100 random seeds for 50 particles and 5000 steps. The proximal map for all algorithms is computed approximately.